1. Linearize the following equations: (a) f(x) = x1/3 + cos(x) around x, = 7/2. (b) g(x) = -3x3 + sin(4x2) around xo = 2.
Answer #1

cat do= II € The given equation is f(0) = x3 + cos(8) defined of the funtion The lineasize is given by by L(x) <(x) = f(x) +C#) - 0) (+)3 -1 is given by the funchon linearizah on až do = 20= f(Q0) + f(20) (%-no) - H D + f(I) (x-1). - 9 % +{{*)(!)--The ginen function at no=2 f(2)= - 3n3 + sin(4x) given by linearization the function is 2(8) = f(M) + flexo) (X-Xo) f(x0) isf(2)= -24 + 0.275 = -23.725 1 (2) = -20.64 by linearization the Equation is given 2(X) = f(90) + f( The) (-dok) = –23.725 +